OpenStudy (anonymous):

Find the intersections of y=x+2 and y^2=4x

OpenStudy (anonymous):

sub x=y-2 into y^2=4x

OpenStudy (dumbcow):

set equations equal to each other,after solving each for y y = x+2 y = sqrt(4x) -> sqrt(4x) = x+2 square both sides 4x = x^2 +4x +4 set equal to 0 x^2+4 = 0 x^2 = -4 x = +-sqrt(-4) = +-2i No real solution

OpenStudy (akshay_budhkar):

nice pic dumbcow

OpenStudy (dumbcow):

heres a graph of both functions so you can see they don't intersect

OpenStudy (dumbcow):

thanks akshay ! Go Cards

OpenStudy (akshay_budhkar):

lol dunbcow.. you could have used wolfram for the graph right?

OpenStudy (dumbcow):

oh yeah..oh well im new to wolfram

OpenStudy (akshay_budhkar):

lol you just type the two equations and done every dam thing comes in front of you.. wait a sec here

OpenStudy (akshay_budhkar):

http://www.wolframalpha.com/input/?i=+y%3Dx%2B2%2C+y^2%3D4x

OpenStudy (akshay_budhkar):

on adding a dx you also get integral and differential of the terms you want

OpenStudy (dumbcow):

pretty nice software

OpenStudy (akshay_budhkar):

lol yea and widely used to

OpenStudy (dumbcow):

i found, graphcalc.com that i use, it will do 3d graphs as well with 2 variables

OpenStudy (akshay_budhkar):

wolfram does that too i believe its a all in 1 package.... I know it i used graphcalc before wolfram but now as i get every thing here i shifted to wolfram

OpenStudy (dumbcow):

ok sweet, i'll prob switch to wolfram then